OpenStudy (anonymous):

A body of mass m, moving at velocity v, collides with a stationary body of the same mass and sticks to it. The question asks for the the momentum and kinetic energy of the two bodies after the collision. My question is why the momentum is mv not 2mv ?

OpenStudy (anonymous):

it depends on what type of collision it is..elastic or inelastic

OpenStudy (anonymous):

But the question doesn't mention anything else.

OpenStudy (kmazraee):

hey goutham! she said that STICK TO IT!

OpenStudy (kmazraee):

because the initial velocity of the stationary mass equals to zero! so : p=(m+m)v/2

OpenStudy (anonymous):

why is v/2 ?

OpenStudy (kmazraee):

m1v1+m2v2=(m1+m2)V so (m1+m2)V=m1v1 because of m1=m2 so V=v1/2 V is the final velocity of body that consist of two mass

OpenStudy (kmazraee):

notice that v2 is the initial velocity of stationary mass

OpenStudy (anonymous):

oh ok i got it thanks

OpenStudy (kmazraee):

ure wellcome!